OpenStudy (anonymous):

Find the length of the curve y= (e^x+1/e^x-1)

OpenStudy (anonymous):

this problems are always cooked up to be easy but you do need an interval right?

OpenStudy (anonymous):

also we need to know if this is \[\frac{e^x+1}{e^x-1}\] which is what i suspect, or something else

OpenStudy (anonymous):

yeah the interval is a<x<b, a>0

OpenStudy (anonymous):

and oops i forgot, there is an ln in front of it

OpenStudy (anonymous):

do me a favour and write it with the equation editor that will help a great deal

OpenStudy (anonymous):

uh i dont think i can use it through my tablet, sorry

OpenStudy (anonymous):

ok let me try \[\ln\left(\frac{e^x+1}{e^x-1}\right)\] how does that look?

OpenStudy (anonymous):

yes thats exactly how it is :)

OpenStudy (anonymous):

ok since i can't compute for beans, lets cheat and find the derivative

OpenStudy (anonymous):

oh actually it is not that hard if we rewrite it as \[\ln(e^x+1)-\ln(e^x-1)\] you get \[\frac{e^x}{e^x+1}-\frac{e^x}{e^x-1}\] pretty fast, but then you have to actually do the subtraction

OpenStudy (anonymous):

so lets cheat and say it is \[-\frac{e^{2x}}{e^{2x}-1}\]

OpenStudy (anonymous):

hmmm but when you take the derivative dont you just get (1/(e^x+1/e^x-1))?

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

you would have to use the chain rule, and it would suck big time, because one you flipped it and got \[\frac{e^x-1}{e^x+1}\] you would still have to multiply by the derivative of \(\frac{e^x+1}{e^x-1}\) which would require the quotient rule

OpenStudy (anonymous):

i.e. it is true that \[\frac{d}{dx}\ln(x)=\frac{1}{x}\] but \[\frac{d}{dx}[f(x)]=\frac{f'(x)}{f(x)}\] you really don't want to do that always best to "simplify" the log first then take the derivative clear? (i hope)

OpenStudy (anonymous):

@satellite73 You made a mistake in simplification It is -2*e^x/(e^2x - 1)

OpenStudy (anonymous):

course we are still not done now we have to square that sucker, which is not too bad, it is \[\frac{4e^{2x}}{(e^{2x}-1)^2}\]

OpenStudy (anonymous):

@bhaskarbabu so i did good catch but the square is right

OpenStudy (anonymous):

then we have to add 1 to that

OpenStudy (anonymous):

when i said it was easy, i meant they are cooked up so that the integral is easy the algebra is a bear

OpenStudy (anonymous):

yeah the square is right @satellite73

OpenStudy (anonymous):

lol thanks god now we have to add 1 and i bet $4 it is a perfect square when we do it (i hope any ways)

OpenStudy (anonymous):

yup i win the $4

OpenStudy (anonymous):

think of all the work some poor underpaid graduate student did to get this problem cooked up so that when you take the derivative square it add 1 you get a perfect square! sheer genius

OpenStudy (anonymous):

think we lost @TheAni-Trix

OpenStudy (anonymous):

lol kinda

OpenStudy (anonymous):

oh no, you are back! ok once all the algebra is done you end up with a perfect square in fact, wolfram tells me it is \(\cosh^2(x)\) althoug i never would have recognized that

OpenStudy (anonymous):

back up like 3 steps

OpenStudy (anonymous):

ok we can back up to wherever you like

OpenStudy (anonymous):

which step?

OpenStudy (anonymous):

to the simplification of the e^x/e^x+1-e^X/e^x-1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

spoiler alert, i used wolfram but we can do it without it it is a matter of subtracting \[\frac{e^x}{e^x+1}-\frac{e^x}{e^x-1}=\frac{e^x(e^x-1)-e^x(e^x+1)}{(e^x+1)(e^x-1)}\]

OpenStudy (anonymous):

the numerator "simplifies" to \(-2e^x\) and the denominator multiplies out to \(e^{2x}-1\) giving \[\frac{-2e^x}{e^{2x}-1}\]

OpenStudy (anonymous):

hows that look?

OpenStudy (anonymous):

how is it -2e^x?

OpenStudy (anonymous):

\[e^x(e^x-1)-e^x(e^x+1)\\ =e^{2x}-e^x-e^{2x}-e^x=-2e^x\]

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

okay yeah i got it :) whats next?

OpenStudy (anonymous):

square it that part is easy no wolfram needed you get \[\left(\frac{dy}{dx}\right)^2=\frac{4e^{2x}}{(e^{2x}-1)^2}\]

OpenStudy (anonymous):

ok for the next step?

OpenStudy (anonymous):

hope it is clear what we are aiming for \[\int_a^b\sqrt{1+\left(\frac{dy}{dx}\right)^2}dx\]

OpenStudy (anonymous):

we got \[\left(\frac{dy}{dx}\right)^2=\frac{4e^{2x}}{(e^{2x}-1)^2}\] now we need \[1+\left(\frac{dy}{dx}\right)^2=1+\frac{4e^{2x}}{(e^{2x}-1)^2}\]

OpenStudy (anonymous):

@satellite73 You are awesome!! I can solve this but I cannot explain like you. You are doing agreat job!! keep it up!! :)

OpenStudy (anonymous):

which is \[\frac{(2e^{2x}-1)^2+4e^{2x}}{(e^{2x}-1)^2}\]

OpenStudy (anonymous):

@bhaskarbabu thanks!

OpenStudy (anonymous):

ready for more algebra?

OpenStudy (anonymous):

lol yes so far im understanding

OpenStudy (anonymous):

lol you're back!

OpenStudy (anonymous):

ok when you multiply all that stuff out in the numerator, you are going to get \[e^{4x}+2e^{2x}+1\] which is luckily a perfect square i.e. it is \[(e^{2x}+1)^2\]

OpenStudy (anonymous):

so now \[1+(\frac{dy}{dx})^2=\frac{(e^{2x}+1)^2}{(e^{2x}-1)^2}\]

OpenStudy (anonymous):

that makes \[\sqrt{1+(\frac{dy}{dx})^2}=\frac{e^{2x}+1}{e^{2x}-1}\] as the square root of the square makes it go away

OpenStudy (anonymous):

see cooked up to work out nice

OpenStudy (anonymous):

okay give me a sec to look over all this and see if i understand it all

OpenStudy (anonymous):

okay give me a sec to look over all this and see if i understand it all

OpenStudy (anonymous):

take your time

OpenStudy (anonymous):

almost all of it is algebra, fraction algebra

OpenStudy (anonymous):

okay i got it up to the last point

OpenStudy (anonymous):

are you good to here \[1+(\frac{dy}{dx})^2=\frac{(e^{2x}+1)^2}{(e^{2x}-1)^2}\]?

OpenStudy (anonymous):

yeah lol i understood all of tha

OpenStudy (anonymous):

and then take the square root and get \[\int\sqrt{1+(\frac{d}{dx})^2}=\int \frac{e^{2x}+1}{e^{2x}-1}dx\]

OpenStudy (anonymous):

there are lots of ways to do that integral i will let you do it unless you still need help, but it can't be that hard after all that work

OpenStudy (anonymous):

okay yeah i got that too

OpenStudy (anonymous):

then you are done right?

OpenStudy (anonymous):

um yeah i got how to do it but can you give me the answer so i can see if i did

OpenStudy (anonymous):

it right

OpenStudy (anonymous):

i guess there are several ways, but partial fractions is probably best

OpenStudy (anonymous):

actually i would do this \[\int \frac{e^{2x}+1}{e^{2x}-1}dx=\int \frac{e^{2x}-1+2}{e^{2x}-1}dx=\int 1dx+\int\frac{2}{e^{2x}-1}dx\]

OpenStudy (anonymous):

a neat trick easier algebra than division

OpenStudy (anonymous):

hmmm i see what you did but i dont think i could come up with that on my own

OpenStudy (anonymous):

yeah it is a nice trick to know but maybe it doesn't really help that much in this case of course you do the mental u-sub and think of it as \[\frac{1}{2}\int \frac{e^u+1}{e^u-1}du\]

OpenStudy (anonymous):

actually it is easier than i am making out break in to two pieces \[\int \frac{e^u}{e^u-1}du+\int \frac{1}{e^u-1}du\] then the first one is obviously \[\ln(e^u-1)\] by substitution

OpenStudy (anonymous):

the second one you can make \(z=e^u-1, z+1=e^u, \ln(z+1)=u, \frac{1}{z+1}dz=du\) and solve \[\int \frac{1}{(z+1)z}dz\] using partial fractions

OpenStudy (anonymous):

actually you get the same answer using either method you decide which is easier

OpenStudy (anonymous):

lol yeah im more used to using the other way with u